Grades assigned by an economics instructor

have historically followed a symmetrical distribution:

5% A’s, 25% B’s, 40% C’s, 25% D’s, and

5% F’s. This year, a sample of 150 grades was drawn

and the grades (1 = A, 2 = B, 3 = C, 4 = D, and

5 = F) were recorded. Can you conclude, at the

10% level of significance, that this year’s grades are

distributed differently from grades in the past?


1
Expert's answer
2021-01-11T18:58:03-0500

Grades are assigned by an economics instructor have historically followed a symmetrical distribution.

The claim is that the grades are distributed differently from the grades in the past.

From the given data we have, the null and alternative hypotheses are

H0: p1 = 0.05, p2 = 0.25, p3 = 0.40, p4 = 0.25, p5 = 0.05

H1: At least one pi is not equil to its specific value

The test statistic for the Chi-Square goodness-of-fit (χ2) is given by,

"χ^2 = \sum^{k}_{i=1}\frac{(f_i-e_i)^2}{e_i}"

ei = expected frequencies

fi = observed frequencies

Given level of significance α=0.10

Using MINITAB, the following are the steps to get the χ2-statistic for the Account receivable data.

1) Click Stat, Table, and Chi-square Goodness-of-fit Test (One Variable)….

2) Type the observed values into the Observe Counts and specify the variable name.

3) Click Proportions specified by historical counts and Input constants type the values of the proportions under the null hypothesis (0.05, 0.25, 0.40, 0.25, 0.05)

4) Click OK.

We get the following output



From the above output,

The test statistic χ2 = 14.0667

The P-value is 0.007

Here we observe that, the P-value is less than the level of the significance 0.10, so we reject the null hypothesis. Therefore, there is enough evidence to support the claim that the grades are distributed differently from the grades in the past.